### v1

@@ 5 Evaluation Metrics
Evaluating the performance of DM-powered image augmentation methods is crucial to understanding their impact on visual image analysis tasks. This section outlines the methodologies used to assess both the efficacy and efficiency of such frameworks.
@@ 5.1.1 Improvement in Model Performance
Analyzing these metrics facilitates towards determining the extent to which augmented images enhance the downstream model’s predictive capabilities.
@@ 5.2 Qualitative Evaluation
Qualitative evaluation involves subjective assessments of the visual realism and relevance of the augmented images. Such types of evaluations are crucial for ensuring that the synthetic data generated by DMs is not only technically sound, but also perceptually convincing and contextually appropriate.
@@ 6.1 Computational Cost and Efficiency
A significant factor associated with the application of any DM-based architecture concerns its very high computational cost. In particular, DM require substantial computational resources and can be time-consuming for both training and inference, which may delay their development and deployment. On the contrary, a trained Generative Adversarial Network (GAN) Generator synthesizes an output image directly, i.e., with just one forward pass. The iterative denoising process of DMs renders it relatively difficult to scale such methods to real-world applications, large datasets and complex tasks. For example, SD requires more than 200,000 GPU hours [117] to be trained on the LAION-5B dataset with one A100 40GB GPU. Approaches such as DDIM and DPM-Solver++ [91] target to speed up the sampling process and to improve efficiency. However, further research is needed to develop more efficient architectures and sampling methods that can generate high-quality images with reduced computational overhead.
@@ 6.2 Lack of Fine-grained Control and Interpretability
Interpretability and controllability of DMs is generally problematic, making it challenging to understand how they generate their outputs (e.g., given specific conditions). The users have limited ability to precisely control specific attributes, objects, or regions in the generated images, which can hinder the usefulness of these methods for certain applications. Developing methods for interpreting the output of DM-based models is an important area of research. Approaches such as classifier-free guidance, cross-attention control or spatial conditions (e.g., masks, regions etc.) have been proposed to mitigate these issues and to enable control of the sampling process. However, further research is required in this area.
@@ 6.3 Limited Diversity and Realism of Generated Data
One of the most prevalent challenges is the (relatively) constrained diversity and realism of the synthetic images generated by DMs. While DMs have shown impressive capabilities in generating high-quality images, they often struggle to capture the full diversity and complexity of real-world data distributions. This limitation can lead to a domain gap between the synthetic and real data, which can hinder the effectiveness of using the generated data for training downstream tasks, like image classification and object detection. Several methods have been introduced so far that aim to address this issue by incorporating approaches that improve the diversity and realism of the generated samples, such as using language enhancement and post-filtering (e.g., CLIP-based filtering). However, more intense and in-depth research is needed to develop more advanced methods that can generate truly diverse and realistic synthetic data across various domains.
@@ 6.4 Model Overfitting and Catastrophic Forgetting
One of the most common challenges faced by many of the methods covered in this study comprises the issue of model overfitting and catastrophic forgetting. In particular, overfitting occurs when a model learns to fit the training data too closely, at the expense of generalization to new, unseen data. This can be especially problematic when working with limited training data, as is often the case in domains like medical imaging. On the other hand, catastrophic forgetting refers to the tendency of trainable models to forget previously learned representations, when being finetuned on new data or tasks. This is a major hurdle for approaches that aim to adapt pretrained models to specific domains or styles. Directly finetuning large models on limited datasets can lead to a rapid loss of the originally learned knowledge structures. Recent approaches that target to mitigate this [189, 203, 187] leverage various methodologies, including selective parameter updates, dual-stream architectures, and careful management of the denoising process, to maintain model generalization while adapting to new tasks or concepts.
@@ 6.5 Evaluation Metrics and Benchmarks
Evaluating the effectiveness of DM-powered image augmentation remains a challenge itself. Traditional metrics like FID and KID may not fully assess/capture the quality and diversity of the generated samples, especially for complex or specialized domains. Additionally, the lack of standardized benchmarks and evaluation protocols makes it difficult to compare different methods and to assess their generalization and augmentation abilities. To this end, developing more comprehensive and both general as well as domain-specific evaluation metrics, while also establishing common benchmarks and datasets, comprises an important area for future research for enabling more rigorous and consistent evaluation.
@@ 7 Conclusion
Image augmentation comprises a fundamental task in modern computer vision, since it allows the enhancement of training datasets with realistic synthetic samples, permiting the automatic context- and semantics-aware editing of given reference images, etc. Diffusion Models (DMs) have shown significant promise in generating realistic and diverse images, capturing complex relationships and structures in high-dimensional image data. Moreover, the ability to condition the generation process using class labels, textual descriptions, or visual prompts allows for targeted augmentation, generating images that fulfill specific requirements based on the task at hand.
@@
Acknowledgements The research leading to the results of this paper has received funding from the European Union’s Horizon Europe research and development programme under grant agreement No. 101073876 (Ceasefire).

### v2

@@ 5 Evaluation Metrics
Evaluating the performance of DM-powered image augmentation methods is crucial to understanding their impact on visual image analysis tasks. This section outlines the methodologies used to assess both the efficacy and efficiency of such frameworks, while a detailed quantitative comparison of DM-based and traditional image augmentation methods is also provided.
@@ 5.1.1 Improvement in Model Performance
Analyzing these metrics makes it easier to determine the degree of improvement in the downstream DNN’s predictive capabilities, which is introduced by training set augmentation.
@@ 5.2 Qualitative Evaluation
Qualitative evaluation involves subjective assessments of visual realism and relevance of the augmented images. Such types of evaluations are crucial for ensuring that the synthetic data generated by DMs is not only technically sound, but also perceptually convincing and contextually appropriate.
@@ 5.2.2 Relevance and Contextual Appropriateness
5.3 Quantitative Evaluation of DM-based and Traditional Image Augmentation Methods
Dataset BASIC-AUG Method DM Method BASIC-AUG (%) DM-AUG (%) Gain (%) Classification - Top-1 Accuracy Aircraft [134] Guided-SR [97] [33] 77.38 84.79 +9.57 Cutmix [250] [222] 89.44 90.25 +0.90 CAL-AUG [164] [136] 81.90 87.40 +6.71 GuidedAP [97] [90] 84.32 85.84 +1.80 GuidedAP [97] [89] 84.32 85.76 +1.70 Caltech101 [63] RandAugment [46] [117] 58.55 59.17 +1.05 RandAugment [46] [66] 60.10 85.30 +41.93 RandAugment [46] [262] 57.80 65.10 +12.61 Cars [108] RandAugment [46] [117] 86.55 88.53 +2.28 RandAugment [46] [66] 65.80 87.10 +32.37 RandAugment [46] [262] 43.20 75.70 +75.23 Guided-SR [97] [33] 91.01 93.04 +2.23 Cutmix [250] [222] 94.73 95.21 +0.50 RandAugment [46] + Cutmix [250] [136] 92.70 93.80 +1.18 GuidedAP [97] [90] 90.27 91.30 +1.14 GuidedAP [97] [89] 90.27 91.26 +1.09 CIFAR-100 [109] GuidedAP [97] [89] 81.20 82.50 +1.70 Cutmix [250] [117] 77.56 75.75 -2.37 GuidedMixup [97] [90] 81.20 82.58 +1.69 Flowers [144] RandAugment [46] [117] 41.97 45.61 +8.67 Cutmix [250] [262] 83.80 88.30 +5.36 GuidedMixup [97] [222] 99.40 99.54 +0.14 iWildCam [106] RandAugment [46] [160] 76.78 85.37 +11.18 Cutmix [250] [56] 77.56 84.87 +9.42 CUB [214] Cutmix [250] [160] 70.48 73.16 +3.80 Co-Mixup [102] [33] 79.41 80.82 +1.77 Cutmix [250] [222] 87.23 87.56 +0.37 Pets [148] RandAugment [46] [117] 57.45 73.71 +28.39 RandAugment [46] [66] 61.50 86.50 +40.65 RandAugment [46] [262] 48.00 73.40 +52.91 Few-shot Classification (8 examples per class) - Top-1 Accuracy MS COCO [122] Std aug [207] 42.00 47.00 +11.90 Std aug [124] 47.00 57.00 +21.67 Object Detection - mAP PASCAL VOC [60] RandAugment [46] [120] 78.20 79.10 +1.15 DIOR-R [38] CopyPaste + Flip [58] [204] 38.75 41.69 +7.58 Out-of-Distribution Classification - OOD Accuracy Waterbirds [170] RandAugment [46] [56] 30.32 46.63 +53.79 Mixup [258] [33] 72.52 76.16 +5.03 Cutmix [250] [222] 71.23 72.47 +1.74 Segmentation - mIoU KITTI [65] Cutout [49] [130] 78.65 82.20 +4.51 Cutout [49] [130] 77.65 78.34 +0.88 Cutout [49] [130] 92.35 92.72 +0.40 PASCAL VOC [60] Std aug [181] 78.10 78.90 +1.02 Medical Image Classification - Top-1 Accuracy OrganMNIST [239] RandAugment [46] [262] 79.60 80.70 +1.38 PathMNIST [239] RandAugment [46] [262] 79.20 86.90 +9.72 BreastMNIST [239] RandAugment [46] [262] 68.70 77.40 +12.66 Shenzhen TB [91] RandAugment [46] [66] 75.50 83.50 +10.59
Having presented the various quantitative and qualitative evaluation metrics, this subsection realizes a comprehensive quantitative comparison among DM-based and traditional image augmentation methods. In particular, Table 4 reports the performance merits of traditional (namely the approaches of [97, 250, 164, 46, 102, 58, 258, 49]) and key DM-based augmentation methods, applied separately to common downstream tasks, namely image classification, object detection and semantic segmentation. The table illustrates the reported improvement that each DM-based augmentation method (denoted DM-AUG) induces, compared to the considered traditional one (denoted BASIC-AUG).
From the observation of Table 4, it can be seen that almost all DM-based augmentation methods demonstrate performance improvements across a wide range of tasks and datasets, over conventional augmentation approaches. The improvement level varies across different tasks, exhibiting substantial (or outstanding) gains in some cases and more modest enhancements in others. It needs to be highlighted though that this increase in performance is also associated with correspondingly significant computational overhead for DM-based methods. In particular, DM-powered generation time may vary considerably across methods and hardware setups, ranging between 0.43-6.6 seconds per image using high-performing GPUs (e.g., H800, V100, RTX3090, etc.), while traditional methods (e.g., Cutout, GridMask, etc.) operate in approximately 0.008 seconds per image [262, 66]. Indicatively, the total generation time of DM-based methods for whole datasets may vary for relatively small datasets from 2 hours (e.g., Watebirds [170]) up to 7 hours (e.g., iWildCam [106]), while certain methods may require up to 10 hours to generate 100K images [57, 54]. Apart from computational cost, DM-powered methods also often require additional storage space for maintaining the generated images, unlike traditional augmentation approaches that can be performed on-the-fly [90]. Moreover, prompt quality dependency is usually a common bottleneck that requires proper construction and careful selection; poor prompt selection could lead to unrealistic or unsuitable generated images that may degrade performance [89, 90].
@@ 6.1 Computational Cost and Efficiency
A significant factor associated with the application of any DM-based architecture concerns its very high computational cost. In particular, DM require substantial computational resources and can be time-consuming for both training and inference, which may delay their development and deployment. On the contrary, a trained Generative Adversarial Network (GAN) Generator synthesizes an output image directly, i.e., with just one forward pass. The iterative denoising process of DMs renders it relatively difficult to scale such methods to real-world applications, large datasets and complex tasks. For example, SD requires more than 200,000 GPU hours [165] to be trained on the LAION-5B dataset with one A100 40GB GPU. Approaches such as DDIM and DPM-Solver++ [126] target to speed up the sampling process and to improve efficiency.
Given this context, a significant avenue for future research comprises the development of more efficient DM architectures and optimization methods. Promising directions include knowledge distillation approaches [128] to create lightweight models [195], quantization [184] or advanced parallelization strategies [187]. Recent work on progressive distillation [174] and one-step generation [244] shows potentials for dramatically reducing inference time. Moreover, research into specialized acceleration [131, 221] could facilitate towards bridging the efficiency gap between DMs and other generative approaches.
@@ 6.2 Lack of Fine-grained Control and Interpretability
Interpretability and controllability of DMs is generally problematic, making it challenging to understand how they generate their outputs (e.g., given specific conditions). The users have limited ability to precisely control specific attributes, objects, or regions in the generated images, which can hinder the usefulness of these methods for certain applications.
@@ 6.3 Limited Diversity and Realism of Generated Data
One of the most prevalent challenges is the (relatively) constrained diversity and realism of the synthetic images generated by DMs. While DMs have shown impressive capabilities in generating high-quality images, they often struggle to capture the full diversity and complexity of real-world data distributions. This limitation can lead to a domain gap between the synthetic and real data, which can hinder the effectiveness of using the generated data for training downstream tasks, like image classification and object detection.
Several methods have been introduced so far that aim to address this issue by incorporating approaches that improve the diversity and realism of the generated samples, such as using language enhancement [55] and post-filtering (e.g., CLIP-based filtering [101]). Further relevant research towards the direction of domain adaptation [40] or style transfer [113] may facilitate towards bridging the gap between real and synthetic image distributions.
@@ 6.4 Model Overfitting and Catastrophic Forgetting
One of the most common challenges faced by many of the methods covered in this study comprises the issue of model overfitting and catastrophic forgetting [47]. In particular, overfitting occurs when a model learns to fit the training data too closely, at the expense of generalization to new, unseen data. This can be especially problematic when working with limited training data, as is often the case in domains like medical imaging. On the other hand, catastrophic forgetting refers to the tendency of trainable models to forget previously learned representations, when being finetuned on new data or tasks. This is a major hurdle for approaches that aim to adapt pretrained models to specific domains or styles. Directly finetuning large models on limited datasets can lead to a rapid loss of the originally learned knowledge structures.
Recent approaches that target to mitigate this [253, 270, 251] leverage various methodologies, including selective parameter updates, dual-stream architectures, and careful management of the denoising process, to maintain model generalization while adapting to new tasks or concepts. Future research along the promising directions of relevant continual learning [191] or metalearning [255] approaches may bear the potentials for further alleviating any remaining issues.
@@ 6.5 Evaluation Metrics and Benchmarks
Evaluating the effectiveness of DM-powered image augmentation remains a challenge itself. Traditional metrics like FID and KID may not fully assess/capture the quality and diversity of the generated samples, especially for complex or specialized domains. Additionally, the lack of standardized benchmarks and evaluation protocols makes it difficult to compare different methods and to assess their generalization and augmentation abilities.
To this end, developing more comprehensive and both general as well as domain-specific evaluation metrics, while also establishing common benchmarks and datasets, comprises an important area for future research for enabling more rigorous and consistent evaluation [15].
@@ 6.6 Ethical Considerations and Bias
Such concerns may partially be mitigated by technical measures, which, however, would require research on detecting and mitigating biases in both training and generated images. Similarly, creating fairness-aware training procedures [64] and establishing ethical guidelines for data collection and model deployment [256] are potentially worthy avenues for investigation, towards more transparent and trustworthy DMs.
@@ 7 Conclusion
Image augmentation comprises a fundamental task in modern computer vision, since it allows the enhancement of training datasets with realistic synthetic samples, permitting the automatic context- and semantics-aware editing of given reference images, etc. Diffusion Models (DMs) have shown significant promise in generating realistic and diverse images, capturing complex relationships and structures in high-dimensional image data. Moreover, the ability to condition the generation process using class labels, textual descriptions, or visual prompts allows for targeted augmentation, generating images that fulfill specific requirements based on the task at hand.
@@
Acknowledgments The research leading to the results of this paper has received funding from the European Union’s Horizon Europe research and development programme under grant agreement No. 101073876 (Ceasefire).